OpenStudy (anonymous):

A pharmacist wishes to mix a solution that is 5% Minoxidil. She has on hand 80 ml of a 4% solution and wishes to add some 7% solution to obtain the desired 5% solution. How much 7% solution should she add?

OpenStudy (anonymous):

0.04*80+0.07x=0.05(80+x) solve it for x= your answer

OpenStudy (anonymous):

The 80 ml solution has (.04)(80) ml of the medicine, or 3.3 ml of medicine.The pharmacist adds V ml of 7% solution, or .07V ml of medicine. The total amoune of medicine is (.07V + 3.2) mg. The total volume of solution is (V + 80) ml, so (.07V + 3.2) / (V + 80) = .05. Solve for V.

OpenStudy (anonymous):

I got V = 40 ml.

OpenStudy (anonymous):

me too!

OpenStudy (anonymous):

Looks like 40's the answer:^)